# SkillDiffuser: Interpretable Hierarchical Planning via Skill Abstractions in Diffusion-Based Task Execution

### v1

@@ Abstract
Diffusion models have demonstrated strong potential for robotic trajectory planning. However, generating coherent and long-horizon trajectories from high-level instructions remains challenging, especially for complex tasks requiring multiple sequential skills. We propose SkillDiffuser, an end-to-end hierarchical planning framework integrating interpretable skill learning with conditional diffusion planning to address this problem. At the higher level, the skill abstraction module learns discrete, human-understandable skill representations from visual observations and language instructions. These learned skill embeddings are then used to condition the diffusion model to generate customized latent trajectories aligned with the skills. This allows for generating diverse state trajectories that adhere to the learnable skills. By integrating skill learning with conditional trajectory generation, SkillDiffuser produces coherent behavior following abstract instructions across diverse tasks. Experiments on multi-task robotic manipulation benchmarks like Meta-World and LOReL demonstrate state-of-the-art performance and human-interpretable skill representations from SkillDiffuser. More visualization results and information could be found on our website.
@@ 1 Introduction
Figure 1: Comparison of SkillDiffuser and previous language conditioned diffusers. SkillDiffuser utilizes high-level abstraction to translate visual observations and language instructions into human understandable skills with language grounding. It then enables the low-level diffusion model condition on these skills, not only improving the execution performance of long-horizon, multi-step composite tasks but greatly enhancing the generalization and adaptability of the framework.
@@
Recent research [17, 18, 6, 5] has demonstrated stronger generation abilities than previous generative models that help enhance reinforcement learning across various dimensions, including the generation of action trajectories [18, 1], representation of policies [48, 4], and synthesis of data [13, 22]. However, their ability to generate coherent trajectories for intricate tasks still poses challenges in terms of performance and generalizability, as these tasks require the completion of abstract instructions comprising numerous coordination-intensive sequential steps.
@@
Previous approaches, [13, 1], such as Decision Diffuser, aim to tackle this challenge by decomposing complex tasks into simpler sub-skills, organized within a predefined skill library. These methods rely on conditioning the diffusion model with one-hot skill representations to generate trajectories for each of these sub-skills. Nevertheless, these techniques encounter difficulties in attempting to autonomously learn end-to-end from a wide range of datasets, which hinders their scalability and ability to achieve end-to-end learning. In addition, without explicitly learning reusable skills, models cannot capture intricate inter-step dependencies and constraints, yielding fragmented and illogical trajectories. Decomposing ambitious instructions into learnable sub-goals and skills can better enable models to follow logical step sequences, respect task structures, and transfer common procedures (e.g., reusable skill abstraction and adaptable skill-based diffusion) between different tasks. This skill-centric paradigm paves the way for diffusion models that can interpret and execute elaborate, abstract instructions necessitating numerous sequential steps.
@@
Inspired by the above observations, we propose SkillDiffuser, a hierarchical planning framework unifying high-level skill learning with low-level conditional diffusion-based execution. As in Fig. 1, compared with previous language-conditioned diffusers, SkillDiffuser is able to interpret and execute complex instructions end-to-end with higher transferability. SkillDiffuser induces interpretable sub-latent goals by learning reusable skills tailored to the task instructions. The framework conditions a diffusion model on these learned skills to generate customized, coherent trajectories aligned with the overall objectives. By combining hierarchical skill decomposition with conditional trajectory generation, SkillDiffuser can produce consistent, skill-oriented behavior without the need for a predefined skill library and fully observed states. This end-to-end approach with learnable skills equips SkillDiffuser to follow abstract instructions across diverse tasks.
@@
SkillDiffuser works as follows. 1) A skill predictor with vector quantization [43] is used for high-level skill learning to distill tasks into discrete and interpretable skills. Rather than forecasting skill durations, we adopt a fixed prediction horizon – predicting skills at regular intervals. This horizon-based discretization process seamlessly integrates visual and linguistic inputs into a cohesive skill set guiding the low-level diffusion model. 2) For skill-based trajectory generation, our diffusion model adopts a classifier-free conditioning approach, directly embedding skills into the model without additional classifiers. This allows for generating diverse state trajectories aligned with skill specifications while avoiding overfitting to a closed dataset. 3) To enable action inference from predicted states, we train an inverse dynamics network to decode motions between two consecutive frames generated. By separating state prediction from motion decoding, SkillDiffuser yields a fully adaptive framework for directing diverse embodiments via transferable state-space plans.
@@
We evaluate the model’s performance in both skill learning and multi-task planning on the LOReL [28] Sawyer and Meta-World [49] datasets, with crucial experimental settings by considering robotic agents in real-world scenarios must operate with limited state information, chiefly visual observations. Furthermore, we present visualization results of SkillDiffuser, especially on skill learning to illustrate the model does have the ability to abstract high-level skills that are not only effective but also understandable to humans, bringing us closer to intelligent agents that acquire skills in an intuitive manner.
@@ 2.1 Imitation Learning and Multi-task Learning
Multi-task learning approaches often leverage shared representations to learn a spectrum of tasks simultaneously, enhancing the flexibility and efficiency of the learning process. The Meta-World benchmark [49] assesses multi-task and meta reinforcement learning, highlighting the need for algorithms capable of rapid adaptation. Building on this, the Prompting Decision Transformer [47] showcases few-shot policy generalization using task-specific prompts. Additionally, diffusion models have been explored as planners in multi-task settings [13], showing proficiency in generating diverse behaviors across tasks.
While these works represent significant strides in multi-task learning, the reliance on state-based inputs rather than pure visual inputs is a notable limitation, pointing to an important direction for future research.
@@ 2.2 Skill Discovery and Hierarchical Learning
Skill learning, fundamentally, referring to the process by which robots acquire new abilities or refine existing ones, is gaining increasing attention due to its pivotal role in enabling autonomous systems to adapt to new tasks, improve their performance over time, and interact more naturally with humans and complex environments. Traditionally, this domain was influenced by hand-crafted features and expert demonstrations [31].
With the development of deep learning, Eysenbach et al. [8] and A. Sharma et al. [38] investigated skill discovery in learning methods, achieving policies conditioned on learned latent variables and maintaining consistent skill codes throughout trajectories. In the domain of skill learning through language instructions, LISA [12] stands out by sampling multiple skills per trajectory, integrating language conditioning in a unique manner. Other studies have also explored hierarchical techniques in reinforcement learning [50, 26, 21]. These contributions enrich the field of skill discovery and hierarchical learning, offering diverse perspectives and methodologies.
@@ 2.3 Planning with Diffusion Model
Our approach is inspired by classifier-free diffusion guidance [16], which offers a significant advantage over traditional classifier-guided models [5]. By adopting the classifier-free approach, we can circumvent the challenges associated with training a reward model and $Q$-function, which are particularly cumbersome in many real-world planning scenarios of which the complexity is very high. Recent studies have also extended this direction, which use conditional diffusion models to generate customized trajectories. Decision Diffuser [1] is designed to generate trajectories for predefined skills. However, its limited one-hot skill encoding hinders its ability to autonomously learn from large datasets. This highlights the necessity for diffusion models with dynamic, learnable skill abstractions, facilitating robust multi-task learning without explicit state representations or pre-set skills. However, developing such an interpretable, end-to-end skill learning framework within diffusion models is still a challenge, and addressing these limitations is critical for enabling diffusion-based planners to effectively execute complex instructions through learned skill representations.
@@ 3.2 Classifier-free Diffusion Guidance
However, the classifier-guided diffusion model requires an accurate estimation of guidance gradient $\nabla_{\boldsymbol{\tau}}\log p\left(\boldsymbol{y}(\boldsymbol{\tau})\right)$ based on the trajectory classifier $\boldsymbol{y}(\boldsymbol{\tau})$, which may not be feasible or may introduce a separate, complex dynamic programming procedure to estimate a $Q$-function in the training process.
Then, classifier-free guided models offer an alternative, where guidance gradient is eschewed, and instead trajectories with desirable characteristics, i.e. $\boldsymbol{y}(\boldsymbol{\tau})$, are emphasized without explicitly training a classifier. It augments the trajectory generation process with a guidance signal that amplifies the features of high-reward or optimal characteristics that have been implicitly present in the data. Mathematically, the noise to add during the reverse diffusion process is manipulated as:
$\hat{\epsilon}=\epsilon_{\theta}(\boldsymbol{\tau}^{i},\varnothing,i)+\omega(\epsilon_{\theta}(\boldsymbol{\tau}^{i},\boldsymbol{y},i)-\epsilon_{\theta}(\boldsymbol{\tau}^{i},\varnothing,i)),$ (5)
@@
where $\beta$ is a hyperparameter that controls the probability of dropping specific conditional information $\boldsymbol{y}(\boldsymbol{\tau})$. This dropout technique aids in enhancing sample diversity while maintaining context relevance.
@@ 4 Methodology
4.1 Overview
Building upon the motivations discussed above, we present SkillDiffuser, an advanced methodological framework for robust multi-task learning across varied robots. This dynamic approach leverages the cooperation of skill learning at the higher level and a conditional diffusion model at the lower level. An overall framework is illustrated in Fig. 2. Notably, as we utilize language-based representations for abstracted skills, we make task execution based on diffusion interpretable and understandable to humans.
SkillDiffuser is designed to only require visual and language input for operation, without the need for explicit agent state information or predefined skill libraries. By abstracting complex tasks into skills (sub-goals), SkillDiffuser enables the diffusion model to work in concert with these sub-goals, ensuring adaptability and flexibility. This unique fusion enables the underlying planner to function consistently across different tasks, with only changes occurring in the inverse dynamic model, offering an unprecedented approach to task learning.
@@ 4.2 High Level Interpretable Skill Abstractions
Firstly, as we utilize only images as robot state information, we transform the images into latent space features with a fixed image encoder such as R3M [29] before images are processed by the skill predictor. Mathematically, we denote the image encoder by $\Phi_{im}:\mathcal{I}\rightarrow\mathcal{R}^{I}$, where $\mathcal{I}$ represents the space of input images and $\mathcal{R}^{I}$, the resultant feature space, serves to condense the visual information into a form that is conducive for high-level semantics. Simultaneously, we use a language encoder to pre-process the natural language instructions, which is formalized as $\Phi_{lang}:L\rightarrow\mathcal{R}^{L}$, with $L$ being the space of language instructions and $\mathcal{R}^{L}$ the language feature space. The outputs of both encoders are then fed into the skill predictor, which operates to integrate these two modalities.
@@
$\tilde{\boldsymbol{z}}=f\left(\Phi_{im}(\boldsymbol{i}_{t}),\Phi_{lang}(l)\right),\vspace{-2pt}$ (7)
$\boldsymbol{z}=\mathbf{q}(\tilde{\boldsymbol{z}})=\operatornamewithlimits{arg\,min}_{\boldsymbol{z}^{k}\in\mathcal{C}}\|\tilde{\boldsymbol{z}}-\boldsymbol{z}^{k}\|_{2}.\vspace{-3pt}$ (8)
VQ enforces each learnt skill $\boldsymbol{z}$ to lie in $\mathcal{C}$, which can be thought as learning $K$ prototypes or cluster centers for the language embeddings utilizing $k$-means [23] algorithm. This acts as a bottleneck, restricting the flow of language information and aiding in the learning of discrete skill codes. The back-propagation through the non-differentiable quantization operation is achieved by a straight-through gradient estimator, which simply copies the gradients and enables the model to be trained end-to-end.
We make each selected skill code $\boldsymbol{z}$ maintain across a designated number of time steps, called as horizon, subtly navigates the challenge of various sub-task durations by simply retaining the same skill code across multiple horizons. This strategy not only preserves the flexibility required for diverse task execution but also maintains the consistency of the model’s architecture.
Importantly, the discrete nature of the learned skill codes enhances the interpretability and controllability of the system’s behavior, as each skill code is associated with some human-understandable language phrases. An example is depicted in Fig. 4. Through this method, SkillDiffuser can not only learn to perform tasks based on language instructions but also achieve them in a human-interpretable way, allowing for a deeper understanding and control of the decision-making process of specific embodied agents.
@@ 4.3 Low Level Skill-conditioned Diffusion Planning
As discussed in Sec. 1, although existing methods (e.g. Decision Diffuser [1]) have proposed conditional diffusion models with skill constraints, they can only generate trajectories satisfying pre-defined skills, because the conditioning in such models is achieved by a classifier which determines whether a generated trajectory aligns with a specific skill or not. Therefore, they do not truly achieve a diffusion model that can condition on an arbitrary learned skill.
To overcome these constraints and enable diffusion models to plan over a learned continuous spectrum of skills, we propose an approach that leverages the classifier-free guided diffusion model with the skills embedded.
@@
Moreover, another important feature of our approach is its state-only diffusion characteristic, which eschews the direct generation of actions in diffusion model to facilitate more generalizable state predictions. And as a better alternative, we utilize another MLP, denoted as $\Psi(\cdot)$, to perform inverse dynamics after state generation to infer feasible actions that can achieve transitions between two continuous states. Additionally, we integrate observations from the current frame to provide more detailed information for motion prediction and achieve closed-loop control. This process is mathematically expressed as:
$\boldsymbol{a}_{t}=\Psi([\tilde{\boldsymbol{s}}_{t},\tilde{\boldsymbol{s}}_{t+1}],\boldsymbol{i}_{t})\quad\text{for}\quad t=0,\ldots,T-1,$ (9)
@@
As the resulting action sequence $\{\boldsymbol{a}_{0},\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{T-1}\}$ derived from the generated states, it encapsulates the skills to execute the tasks, which allows for remarkable adaptability across multiple tasks. And when faced with a new task, we need only to replace the inverse dynamics model $\Psi(\cdot)$ specific to the new task’s kinematics, with the architecture and parameters of diffusion model unchanged. Such modularity ensures that the generative capabilities of SkillDiffuser are not task-specific but can be leveraged across a diverse range of tasks with varying dynamics, which mirrors the complex skill interplay and adaptability seen in human learning. A schematic diagram of low level skill-conditioned diffusion planning model is shown in Fig. 3.

### v3

@@ Abstract
Diffusion models have demonstrated strong potential for robotic trajectory planning. However, generating coherent trajectories from high-level instructions remains challenging, especially for long-range composition tasks requiring multiple sequential skills. We propose SkillDiffuser, an end-to-end hierarchical planning framework integrating interpretable skill learning with conditional diffusion planning to address this problem. At the higher level, the skill abstraction module learns discrete, human-understandable skill representations from visual observations and language instructions. These learned skill embeddings are then used to condition the diffusion model to generate customized latent trajectories aligned with the skills. This allows generating diverse state trajectories that adhere to the learnable skills. By integrating skill learning with conditional trajectory generation, SkillDiffuser produces coherent behavior following abstract instructions across diverse tasks. Experiments on multi-task robotic manipulation benchmarks like Meta-World and LOReL demonstrate state-of-the-art performance and human-interpretable skill representations from SkillDiffuser. More visualization results and information could be found on our website.
@@ 1 Introduction
Figure 1: Comparison of SkillDiffuser and previous language conditioned diffusers. SkillDiffuser utilizes high-level abstraction to translate visual observations and language instructions into human understandable skills with language grounding. It then enables the low-level diffusion model condition on these skills, not only improving the execution performance of multi-step composition tasks but greatly enhancing the generalization and adaptability of the framework.
@@
Recent research [18, 19, 7, 6] has demonstrated diffusion models’ superior generative capabilities compared to previous models that help enhance reinforcement learning across various dimensions, including the generation of action trajectories [19, 2], policy representation [50, 5], and data synthesis [14, 23]. However, their ability to generate coherent trajectories for intricate tasks still poses challenges in terms of performance and generalizability, as these tasks often require the fulfillment of abstract instructions that consist of numerous coordination-intensive sequential steps.
@@
Previous approaches [14, 2], such as Decision Diffuser, aim to tackle this challenge by decomposing complex tasks into simpler sub-skills, organized within a predefined skill library. These methods rely on conditioning the diffusion model with one-hot skill representations to generate trajectories for each of these sub-skills. Nevertheless, these techniques encounter difficulties in attempting to autonomously learn end-to-end from a wide range of datasets, which hinders their scalability and ability to achieve end-to-end learning [1]. In addition, without explicitly learning reusable skills, models cannot capture intricate inter-step dependencies and constraints, yielding fragmented and illogical trajectories. Decomposing ambiguous instructions into learnable sub-goals and skills can better enable models to follow logical step sequences, respect task structures, and transfer common procedures (e.g. reusable skill abstraction and adaptable skill-based diffusion) between different tasks. This skill-centric paradigm paves the way for diffusion models that can interpret and execute elaborate, abstract instructions necessitating numerous sequential steps.
@@
Inspired by the above observations, we propose SkillDiffuser, a hierarchical planning framework unifying high-level skill learning with low-level conditional diffusion-based execution. As shown in Fig. 1, compared with previous language-conditioned diffusion policies, SkillDiffuser is able to interpret and execute complex instructions end-to-end with higher transferability. SkillDiffuser induces interpretable sub-latent goals by learning reusable skills tailored to the task instructions. The framework conditions a diffusion model on these learned skills to generate customized, coherent trajectories aligned with the overall objectives. By integrating hierarchical skill decomposition with conditional trajectory generation, SkillDiffuser achieves consistent, skill-oriented behavior without relying on a predefined skill library. Moreover, SkillDiffuser is designed to operate solely on visual observations, eliminating the need for robot proprioception (i.e. fully observed states). This end-to-end methodology, featuring learnable skills, enables SkillDiffuser to execute abstract instructions across a variety of tasks efficiently.
@@
SkillDiffuser works as follows. 1) A skill predictor with vector quantization [45] is used for high-level skill learning to distill tasks into discrete and interpretable skills. Rather than forecasting skill durations, we adopt a fixed prediction horizon – predicting skills at regular intervals. This horizon-based discretization process seamlessly integrates visual and linguistic inputs into a cohesive skill set guiding the low-level diffusion model. 2) For skill-based trajectory generation, we utilize a classifier-free diffusion model as policy, with skills directly embedded as guidance. This setting allows for generating multi-modal state trajectories aligned with skill specifications while avoiding overfitting to a closed dataset. 3) To enable action inference from predicted states, we train an inverse dynamics network to decode motions between two consecutive frames generated. By separating state prediction from motion decoding, SkillDiffuser yields a fully adaptive framework for directing diverse embodiments via transferable state-space plans.
@@
We evaluate the model’s performance in both skill learning and multi-task planning on the LOReL [29] Sawyer and Meta-World [51] datasets, with crucial experimental settings by considering robotic agents in real-world scenarios must operate with limited state information, chiefly visual observations. Furthermore, we present success rates on unseen compositional tasks, the reusability and visualizations of learned skills to illustrate the model does have the ability to abstract high-level skills that are not only effective but also understandable to humans, bringing us closer to intelligent agents that acquire skills in a direct manner.
@@ 2.1 Imitation Learning and Multi-task Learning
Multi-task learning approaches often leverage shared representations to learn a spectrum of tasks simultaneously, enhancing the flexibility and efficiency of the learning process. The Meta-World benchmark [51] assesses multi-task and meta reinforcement learning, highlighting the need for algorithms capable of rapid adaptation. Building on this, the Prompting Decision Transformer [49] showcases few-shot policy generalization using task-specific prompts. And diffusion policy has also been explored in multi-task settings [14], which shows proficiency in generating diverse behaviors across tasks. However, unlike methods that leverage state inputs [14, 49] or access robot proprioception [30], SkillDiffuser uses raw visual inputs only.
@@ 2.2 Skill Discovery and Hierarchical Learning
Skill learning, the process by which robots acquire new abilities or refine existing ones, is gaining increasing attention due to its pivotal role in enabling autonomous systems to adapt to new tasks, improve the performance over time, and interact naturally with humans and complex environments. Traditionally, this domain was influenced by hand-crafted features and expert demonstrations [32].
With the development of deep learning, Eysenbach et al. [9] and A. Sharma et al. [40] investigated skill discovery in learning methods, achieving policies conditioned on learned latent variables and maintaining consistent skill codes throughout trajectories. In the domain of skill learning through language instructions, LISA [13] stands out by sampling multiple skills per trajectory, integrating language conditioning in a unique manner.
Our SkillDiffuser follows this way to extract sub-skills of each instruction at the higher-level. But differently, SkillDiffuser employs an adaptable diffusion policy at the lower-level to condition on different sub-skills to generate different actions, which formulates a creative hierarchical planning framework also advancing research on hierarchical techniques of reinforcement learning [52, 27, 22].
@@ 2.3 Planning with Diffusion Model
Our approach is inspired by classifier-free diffusion guidance [17], which offers a significant advantage over classifier-guided models [6]. By adopting the classifier-free approach, we can circumvent the challenges associated with training a reward model and $Q$-function, which are particularly cumbersome in many real-world planning scenarios of which the complexity is very high. Recent studies have also extended this direction, which use conditional diffusion models to generate customized trajectories. Decision Diffuser [2] is an example which is designed to generate trajectories for a predefined skill library. However, unlike our method, it can’t autonomously learn skill abstractions in an end-to-end fashion, which makes it difficult to scale to more tasks. This highlights the necessity for diffusion models with dynamic, learnable skill abstractions, facilitating complex instruction execution.
@@ 3.2 Classifier-free Diffusion Guidance
However, the classifier-guided diffusion model requires an accurate estimation of guidance gradient based on the trajectory classifier $\boldsymbol{y}(\boldsymbol{\tau})$, which may not be feasible and need to introduce a separate dynamic programming procedure to estimate a $Q$-function in the training process.
Thus, classifier-free guidance offer an alternative, which augments the trajectory generation process with a guidance signal that amplifies the features of high-reward or optimal characteristics, i.e. $\boldsymbol{y}(\boldsymbol{\tau})$, that are implicitly present in the data. Mathematically, the noise to add during the reverse denoising process is:
$\hat{\epsilon}=\epsilon_{\theta}(\boldsymbol{\tau}^{i},\varnothing,i)+\omega(\epsilon_{\theta}(\boldsymbol{\tau}^{i},\boldsymbol{y},i)-\epsilon_{\theta}(\boldsymbol{\tau}^{i},\varnothing,i)),\vspace{-5pt}$ (5)
@@
where $\beta$ is a hyperparameter that controls the probability of dropping specific condition $\boldsymbol{y}(\boldsymbol{\tau})$, enhancing sample diversity while maintaining context relevance.
@@ 4 Methodology
4.1 Overview
Building upon the motivations discussed above, we present SkillDiffuser, an advanced methodological framework for robust multi-task learning across various robots. This dynamic approach leverages the cooperation of skill learning at the higher level and a conditional diffusion model at the lower level. An overall framework is illustrated in Fig. 2. Notably, we leverage language-grounded representations to abstract skills, thereby rendering the execution of tasks through our diffusion policy both interpretable and comprehensible to humans.
@@ 4.2 High Level Interpretable Skill Abstractions
Firstly, as we utilize only images as robot state information, we transform the images into latent space features with a fixed image encoder (e.g. R3M [30]). For convenience, we denote the image encoder as $\Phi_{im}:\mathcal{I}\rightarrow\mathcal{R}^{I}$, where $\mathcal{I}$ represents the space of input images and $\mathcal{R}^{I}$, the resultant feature space, serves to condense the visual information into a form that is conducive for high-level semantics. Simultaneously, we use a language encoder to pre-process the natural language instructions, which is formalized as $\Phi_{lang}:L\rightarrow\mathcal{R}^{L}$, with $L$ being the space of language instructions and $\mathcal{R}^{L}$ the language feature space. The outputs of both encoders are then fed into the skill predictor, which operates to integrate these two modalities.
@@
$\tilde{\boldsymbol{z}}=f\left(\Phi_{im}(\boldsymbol{i}_{t}),\Phi_{lang}(l)\right),\vspace{-12pt}$ (7)
$\boldsymbol{z}=\mathbf{q}(\tilde{\boldsymbol{z}})=\operatornamewithlimits{arg\,min}_{\boldsymbol{z}^{k}\in\mathcal{C}}\|\tilde{\boldsymbol{z}}-\boldsymbol{z}^{k}\|_{2}.\vspace{-2pt}$ (8)
VQ enforces each learnt skill $\boldsymbol{z}$ to lie in $\mathcal{C}$, which is equal to learning $K$ prototypes for the language embeddings utilizing $k$-means [24] algorithm. This acts as a bottleneck, restricting the flow of language information and aiding in the learning of discrete skill codes. The back-propagation through the non-differentiable quantization is achieved by a straight-through gradient estimator, which simply copies the gradients and enables model to be trained end-to-end.
In our approach, we apply a consistent skill code $\boldsymbol{z}$ across a defined number of time steps, termed the horizon. This consistent application across multiple horizons adeptly addresses the challenge of varying sub-task durations without altering the horizon itself. Consequently, this strategy not only preserves the flexibility required for diverse task execution but also maintains the model’s architectural stability by avoiding horizon-induced structural changes.
Importantly, the discrete nature of the learned skill codes enhances the interpretability and controllability of the system’s behavior, as each skill code is associated with some human-understandable language phrases. An example is depicted in Fig 6. Through this method, SkillDiffuser can not only learn to perform tasks based on language instructions but also achieve them in a human-interpretable way, allowing for a deeper understanding and control of the decision-making process of specific embodied agents.
@@ 4.3 Low Level Skill-conditioned Diffusion Planning
As highlighted in Section 1, while existing approaches like Decision Diffuser [2] have introduced conditional diffusion models constrained by skills, their capability is limited to generating trajectories that meet only predefined skills. Consequently, these models fall short of realizing a diffusion model capable of conditioning on an arbitrary learned skill. To overcome these constraints and enable diffusion models to plan over a learned continuous spectrum of skills, we propose an approach that leverages the classifier-free guided diffusion model with the skills embedded.
@@
Moreover, following previous works [2, 23, 7], we adopt a state-only diffusion model, which eschews the direct generation of actions in diffusion model. And instead, we utilize another MLP, denoted as $\Psi(\cdot)$, to perform inverse dynamics after state generation to infer feasible actions that can achieve transitions between two continuous states. Additionally, we integrate observations from the current frame to provide more detailed information for motion prediction and achieve closed-loop control. Mathematically, it is:
$\boldsymbol{a}_{t}=\Psi([\tilde{\boldsymbol{s}}_{t},\tilde{\boldsymbol{s}}_{t+1}],\boldsymbol{i}_{t})\quad\text{for}\quad t=0,\ldots,T-1,\vspace{-5pt}$ (9)
@@
As the resulting action sequence $\{\boldsymbol{a}_{0},\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{T-1}\}$ derived from the generated states, it encapsulates the skills to execute the tasks, which allows for remarkable adaptability across multiple tasks. And when faced with a new task, we are only required to change the inverse dynamics model $\Psi(\cdot)$ specific to the new task’s kinematics, with the architecture and parameters of diffusion model unchanged. Such modularity ensures the generative capabilities of SkillDiffuser are not task-specific but can be leveraged across a diverse range of tasks with varying dynamics. A schematic diagram of the low level module is shown in Fig. 3.